OpenStudy (vera_ewing):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the second equation is substituted into the first equation. 2x – 3y = –29 x + 4y = 13 (5 points) 2(4y + 13) – 3y = –29 2(–4y + 13) – 3y = –29 2x – 3(4y + 13) = –29 2x – 3(–4y + 13) = –29

OpenStudy (anonymous):

So first, we need to solve for \(x\) in the second equation.

OpenStudy (anonymous):

\[ \underbrace{x + 4y}_{-4y} = \underbrace{13}_{-4y}\\ x = 13-4y \]

OpenStudy (jdoe0001):

x + 4y = 13 <---- solve for "x" first, as wio pointed out

OpenStudy (vera_ewing):

do we divide by 4 on both sides now?

OpenStudy (anonymous):

No, the \(x\) is isolated so it is good.

OpenStudy (vera_ewing):

so now what? what would the answer be? i think it's B

OpenStudy (anonymous):

You substitute right side into \(x\). The answer is B